OpenStudy (anonymous):

1. Translate the phrase “nine more than two times a number” into an algebraic expression. (1 point) 9x + 2 2x + 9 9x – 2 2x – 9 2. Evaluate the expression 3m – 5 when m = 7. (1 point) 2 16 21 28 3. Simplify 4x + 8y – 3x + 2 y. (1 point) 12x – y 11x + y 6x + 5y x + 10y 4. Factor 24p + 64. (1 point) 24(p + 40) 12(2p + 5) 8(3p + 8) 3(8p + 21) For problems 5 to 12, use Properties of Equality to solve the equation. 5. m + 8 = 22 (1 point) −30 14 −14 30 6. w – 8 = −12 (1 point) −20 −4 4 20 7. 3q = 12 (1 point) 4 9 15 36

OpenStudy (anonymous):

plz help i have more of these

OpenStudy (anonymous):

number 1. is B 2.16 3.D 4.C 5. B 6.-4 7.4

OpenStudy (anonymous):

i have more

OpenStudy (anonymous):

SHOO

OpenStudy (anonymous):

16. z – 4 > 9 (1 point) z > −13 z > −5 z > 5 z > 13 17. q + 7 ≤ − 15 (1 point) q ≤ −22 q ≤ − 8 q ≤ 8 q ≤ 15 18. −3y > 9 (1 point) y < 3 y > 3 y < −3 y > −3 19. ≥ –4 (1 point) k ≥ −20 k ≤ −20 k ≥ 20 k ≤ 20 20. 4g – 8 < 20 (1 point) g > 3 g < 3 g > 7 g < 7 21. + 12 ≤ 8 (1 point) v ≥ 8 v ≤ 8 v ≥ −8 v ≤ −8

OpenStudy (anonymous):

Brandon is bringing bottled water to a picnic. He has 4 full cases of bottles of water, and a bag with 8 more bottles. He has a total of 104 bottles of water. Write and solve an equation to find the number of bottles in each case. (1 point) + 8 = 104; x = 384 – 8 = 104; x = 448 4x + 8 = 104; x = 24 4x – 8 = 104; x = 28 14. Which is a solution of h > −6? (1 point) –10 –8 –6 –4

OpenStudy (anonymous):

plz help me

OpenStudy (anonymous):

hold up

OpenStudy (anonymous):

q/6=24 4 8 30 144 9. 3b – 9 = 15 (1 point) 8 6 2 –2 10. + 10 = –15 (1 point) –125 –25 –5 –1 11. −3(4 − a) = 6 (1 point) −6 −2 2 6 12. 5(x + 6) = 45 (1 point) 15 3 –3 –

OpenStudy (anonymous):

and thats it

OpenStudy (anonymous):

D B C #19 AND 21 ARE MISSING #s 20.D

OpenStudy (anonymous):

hold up

OpenStudy (anonymous):

not sure on the word problem

OpenStudy (anonymous):

no word prblom

OpenStudy (anonymous):

do the answer in order

OpenStudy (anonymous):

plz

OpenStudy (anonymous):

dont know the one under #21

OpenStudy (anonymous):

do it like 123456789 ok the answers

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

8.D 9.A 10 MISSING INFO 11.D 12.C

OpenStudy (anonymous):

ok 10 missing info

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

dont answer 10

OpenStudy (anonymous):

Brandon is bringing bottled water to a picnic. He has 4 full cases of bottles of water, and a bag with 8 more bottles. He has a total of 104 bottles of water. Write and solve an equation to find the number of bottles in each case. (1 point) + 8 = 104; x = 384 – 8 = 104; x = 448 4x + 8 = 104; x = 24 4x – 8 = 104; x = 2 Dont know this one 14. Which is a solution of h > −6? (1 point) answer -4 –10 –8 –6 –4

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

could i get a medal

OpenStudy (anonymous):

hold up

OpenStudy (anonymous):

what about the rest

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

the 21 Qs

OpenStudy (anonymous):

then i will give u the medal

OpenStudy (anonymous):

number 1. is B 2.16 3.D 4.C 5. B 6.-4 7.4 8.D 9.A 10 MISSING INFO 11.D 12.C 59 Carter15 0Brandon is bringing bottled water to a picnic. He has 4 full cases of bottles of water, and a bag with 8 more bottles. He has a total of 104 bottles of water. Write and solve an equation to find the number of bottles in each case. (1 point) + 8 = 104; x = 384 – 8 = 104; x = 448 4x + 8 = 104; x = 24 4x – 8 = 104; x = 2 Dont know this one 14. Which is a solution of h > −6? (1 point) answer -4 –10 –8 –6 –4 15.D 16. B 17 C 18.D 19 MISSING info 20.D 21.missing info

OpenStudy (anonymous):

thats it

OpenStudy (anonymous):

im pressing submit

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

13/21

OpenStudy (anonymous):

thats bad

OpenStudy (anonymous):

your wellcome

OpenStudy (anonymous):

goodbye